OpenStudy (miss.rose):

Drawing Triangles Question 1:

OpenStudy (yaihlyn):

whats the question?

OpenStudy (miss.rose):

@F_Jayyy

OpenStudy (miss.rose):

Which collection of three side lengths can be used to form a triangle? 75 cm, 35 cm, 40 cm 40 cm, 65 cm, 30 cm 40 cm, 30 cm, 75 cm 80 cm, 35 cm, 40 cm

OpenStudy (miss.rose):

@F_Jayyy

OpenStudy (f_jayyy):

well here is a simple tip....the total area of a triangle is 180 :)

OpenStudy (miss.rose):

75+35+40 = 150 40+65+30= 135 40+30+75= 145 80+35+40=155

OpenStudy (f_jayyy):

yup....thats where i get confused....does it give you a choice to choose none?

OpenStudy (miss.rose):

no

OpenStudy (miss.rose):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

@F_Jayyy remember the degrees of a triangles angles should be 180 not the lengths

OpenStudy (f_jayyy):

true true true

rebeccaxhawaii (rebeccaxhawaii):

http://prntscr.com/aml4xs

OpenStudy (miss.rose):

I still don't get it.

rebeccaxhawaii (rebeccaxhawaii):

is this similar to what your lesson has taught you?

OpenStudy (miss.rose):

It's similar, I just don't understand it.

rebeccaxhawaii (rebeccaxhawaii):

for for a, b, and c plug your numbers in

OpenStudy (miss.rose):

a-75 b-35 c-40

rebeccaxhawaii (rebeccaxhawaii):

okay plug it in to what i just gave you

OpenStudy (miss.rose):

I don't get how to do that.

OpenStudy (miss.rose):

a+b>c a-75 b-35 c-40 75+35>40

rebeccaxhawaii (rebeccaxhawaii):

\[a+b>c\] \[a+c>b\] \[b+c>a\] plug your number in

rebeccaxhawaii (rebeccaxhawaii):

numbers*

OpenStudy (miss.rose):

75+35>40 75+40>35 35+40>75

OpenStudy (miss.rose):

@rebeccaxhawaii where did you go?

rebeccaxhawaii (rebeccaxhawaii):

okay now add the numbers

OpenStudy (miss.rose):

what does > mean?

rebeccaxhawaii (rebeccaxhawaii):

> means greater than < means less then so 7 is greater than 3 or 7 > 3

OpenStudy (miss.rose):

110>40 115>40 75>75

OpenStudy (miss.rose):

opps, the second one is supposed to be 115>30

OpenStudy (n8kd_da_beast):

what question is needed?

OpenStudy (miss.rose):

What do you mean? Both the questions are the same. I am just not sure what to do after this point, to get the answer, and I have been in this test for an hour and I haven't even done the first question.

OpenStudy (n8kd_da_beast):

ok . one moment please....

OpenStudy (n8kd_da_beast):

Ok, I really hate disappointing people but I just had a total brain fart and unfortunately cant answer this problem. Sorry :(

OpenStudy (phi):

The easiest way to do these is pick out the biggest number out of the 3 numbers. then make sure the other 2 add up to a *bigger* (not just equal) number.

OpenStudy (miss.rose):

What do you mean?

OpenStudy (miss.rose):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

just plug in your numbers into that formula

OpenStudy (miss.rose):

how?

rebeccaxhawaii (rebeccaxhawaii):

plug a for a plug in b for b and plug in c for c

OpenStudy (miss.rose):

110>40 115>35 75>75

rebeccaxhawaii (rebeccaxhawaii):

is 110 greater than 40? is 115 greater that 35? is 75 greater than 75?

OpenStudy (miss.rose):

yes yes equal

rebeccaxhawaii (rebeccaxhawaii):

so thats not the choice because 75 is not greater than 75 so just do that for each option and you will eventually find one that fits good luck

OpenStudy (miss.rose):

Okay, thanks!

OpenStudy (miss.rose):

@rebeccaxhawaii So am I looking for all the answers to be yes? Or equal?

rebeccaxhawaii (rebeccaxhawaii):

no > means greater than not equal = means equal

OpenStudy (miss.rose):

No what I mean is when you asked me if 110 was greater than 40, and I said yes, when I have the greater formula am I looking for it to be equal or no?

OpenStudy (miss.rose):

@rebeccaxhawaii